# Generalizing CLIP to Unseen Domain via Text-Guided Diverse Novel Feature Synthesis

### v1

@@ Abstract
Vision-language foundation models like CLIP have shown impressive zero-shot generalization, but finetuning on downstream datasets can cause overfitting and loss of its generalization ability on unseen domains. Although collecting additional data from new domains of interest is possible, this method is often impractical due to the challenges in obtaining annotated data. To address this, we propose a plug-and-play feature augmentation method called LDFS (Language-Guided Diverse Feature Synthesis) to synthesize new domain features and improve existing CLIP fine-tuning strategies. LDFS has three main contributions: 1) To synthesize novel domain features and promote diversity, we propose an instance-conditional feature augmentation strategy based on a text-guided feature augmentation loss. 2) To maintain feature quality after augmenting, we introduce a pairwise regularizer to preserve augmented feature coherence within the CLIP feature space. 3) We propose to use stochastic text feature augmentation to reduce the modality gap and further facilitate the process of text-guided feature synthesis. Extensive experiments show LDFS’s superiority in improving CLIP’s generalization ability on unseen domains without collecting data from those domains. The code will be made publicly available.
Abstract
In this supplementary material, we provide more implementation details, additional experiments, and additional ablation studies.
Keywords:
Vision and language Generalization Data augmentation
1 Introduction
@@
In order to further adapt CLIP to downstream tasks, some works propose to fine-tune CLIP via either linear probing or learning optimal soft prompts, such as CoOp [41], CoCoOp [40], MaPle [14], and PromptSRC [15], in a process called prompt learning for downstream datasets. Although these fine-tuning methods can improve CLIP’s performance on downstream datasets, a significant issue is that they can also cause the model to overfit to these datasets, resulting in even worse performance than the original zero-shot CLIP on some related datasets with domain-shift, limiting real-world applicability. One method is to collect additional data from the target domains of interest. However, this approach is often impractical due to the challenges in obtaining data from new or rare domains, further compounded by the costs of annotation for effective adaptation.
To address this issue, we aim to develop a feature augmentation method that can synthesize visual features from novel domains by leveraging the knowledge existing in pre-trained CLIP. Our motivation is based on the observation that the domain-shift direction in CLIP’s text embedding can guide image feature augmentation in CLIP’s image embedding, without requiring access to any images from the target domain [17, 5, 8]. However, directly manipulating the text latent space of CLIP to synthesize either novel domain features (e.g., LADS [5]) or images (e.g., CLIPStyler [17]) cannot effectively improve CLIP’s performance on unseen domains. We attribute this to the tendency of these methods to synthesize visual features that lack diversity within the same class (e.g., LADS in Fig. 4) or generate low-quality images with artifacts (e.g., CLIPStyler in Fig. 12).
@@
In this paper, we propose LDFS (Language-Guided Diverse Feature Synthesis), which synthesizes diverse and high-quality novel domain visual features given source domain training data along with text descriptions of the training data and the target domain of interest only, as illustrated in the left of Fig. 1. We achieve high diversity and high-quality synthesized features using three approaches. 1)Promoting feature diversity with instance-conditional augmentation: We notice that existing text-guided feature augmentation and styler transfer methods [5, 8, 17] use a shared, global augmentation direction for all images in each category, leading to homogeneous synthesized features or images. We address this by introducing a novel adaptation loss that utilizes more nuanced, instance-conditional text descriptions instead of uniform descriptions. This enables our method to synthesize more diverse visual features. As illustrated in Fig. 3, rather than applying a global direction from “a photo of a horse” to “a sketch of a horse” for all horse images, our method employs customized descriptions like “a (sketch) photo of a horse with its head raised” or “a (sketch) photo of a horse grazing on a green field” for individual images. This strategy ensures that the augmentation process is informed by the unique attributes of each image, enhancing feature diversity and capturing richer visual semantics. 2) Preserving feature coherence in adapted space: We observe that the feature augmentation process could inadvertently drift synthesized features from the CLIP sphere, negatively impacting the synthesis quality and losing the class information. This occurs because the CLIP-based augmentation loss overlooks the preservation of pairwise cosine distances for adapted features. To mitigate it, we propose a regularizer to explicitly maintain the pairwise relationship for augmented features so that augmented features lie in the CLIP sphere, achieving better synthesis quality. 3) Reducing the cross-modality representation gap: Despite CLIP’s ability to represent text and image features in a joint space, there persists a modality representation gap, as documented in recent studies [20, 11]. We notice that this modality gap weakens the integration of cross-modality features and thus hampers the adaptation process. To address it, we propose adjusting the text representation closer to its image counterpart by introducing controlled perturbations.
@@ 2 Related Work
Text-guided Image Synthesis. Text-guided image synthesis [8, 17, 25, 16] aims to modify source images through text descriptions. For example, StyleGAN-NADA [8] achieves zero-shot image synthesis by leveraging text prompts, identifying the domain shift in the text space as a directional to guide image generation. Similarly, CLIPStyler [17] finetunes a style transfer network utilizing CLIP embeddings for text-driven style transfer. While these works make progress in generative tasks, they are not designed for tackling domain shift problems and necessitate additional generative models. Moreover, their efficacy is frequently constrained by the quality of the generation process, which often introduce artifacts or wrong classes for downstream adaptation (see Fig. 12). Similarly, LADS [5] and PODA are introduced, proposing to directly manipulate CLIP embeddings to synthesize features for model adaptation. Nevertheless, a common issue among all the above methods is their reliance on a global adaptation direction within the CLIP space, which limits the diversity of synthesized images or features. This sometimes leads to even worse performance than zero-shot CLIP (e.g., LADS compared to CLIP (ZS) in Table 1 to 4 and CLIPStyler compared to CLIP (ZS) in Table 4.2. In contrast to prior efforts, we introduce an instance-conditional local direction adaptation framework to promote both diverse and high-quality visual feature synthesis for effective downstream adaptation.
@@
Model Generalization. Domain adaptation (DA) aims to bridge the distribution gap between the source and target domains so that the model can generalize in target domains. In unsupervised domain adaptation, this is typically achieved by leveraging additional unlabeled data from the target domain and utilizing techniques such as adversarial training [9, 12, 7], and moment matching [27, 21, 3], to minimize distribution disparities, all while reducing the necessity for costly annotations in the target domain. More recently, few-shot domain adaptation [23, 36] has been introduced, allowing models to adapt to a new domain with only a few examples, further relaxing the data requirement. Diverging from DA, our method does not require any data from target domains. Instead, it utilizes descriptive text about the source and target domains, guiding the adaptation process of source images via CLIP [29] models. It’s noteworthy that, while domain generalization [39] also does not require target domain data, it typically requires training on multiple source domain data. Although some methods [33, 28] only require a single source domain data (called single domain generalization), they often lead to limited improvements in data with domain shift. This is in contrast to our method, which is trained with just a single source domain and significantly enhances CLIP’s fintuning performance.
3 Method
@@
In this section, we present our framework based on an instance-conditional local adaptation strategy, a pair-wise regularizer, and text feature augmentation, which synthesizes diverse and high-quality target domain features rather than using original features only. To formulate the task, consider a training set $\mathcal{D}_{train}=\{(x_{s},y_{s})\}_{s=1}^{S}$ where $x_{s}$ is an input image from source domain, $y_{s}$ is the corresponding ground truth, and S is the size of the source training dataset. The task also includes the description $t_{s}$ of the source training domain (e.g.,“a real photo of a {}”) plus the class name (e.g.,horse), and a collection of $k$ descriptions $\{t^{i}_{t}\}_{i=1}^{k}$ of the $k$ unseen target domains $\{\mathcal{D}^{i}_{unseen}\}_{i=1}^{k}$ that cover the same classes as the source dataset. The goal is to adapt the model to work well on the $k$ unseen target domains during testing time.
@@
Our two-stage method, shown in Fig. 2, begins with extracting source image features $f=I(x_{s})$ using CLIP’s frozen image encoder. Then, we enrich both source $t_{s}$ and target text descriptions $t_{t}$ with attributes from a captioning model. The refined descriptions are encoded by CLIP’s text encoder into $t^{\prime}_{s}$ and $t^{\prime}_{t}$. To bridge the modality gap, we apply stochastic, data-independent noise to span the space of text features. A mapper network then maps the source image features $f$ into diverse target features $f^{\prime}$. The mapping process is guided by the domain gap direction between $t^{\prime}_{s}$ and $t^{\prime}_{t}$ using a local adaptation loss $\mathcal{L}_{ld}$. We also employ a sphere regularizer to constrain the synthetic features $f^{\prime}$ within the CLIP sphere. In stage 2, we combine the source image features $f$ with synthetic features $f^{\prime}$ and finetune them to enhance the model performance in unseen domains.
3.1 Text-Guided Diverse Feature Synthesis
@@
Instance-conditional local adaptation loss: In the generative model community, existing text-guided image synthesis methods use CLIP embeddings to guide additional generative models like StyleGAN to translate images through losses like global CLIP loss [25] or directional CLIP loss [8, 5]. However, existing CLIP losses utilize a fixed global direction for all images in a class, limiting the diversity of synthetic images, as illustrated in Fig. 3 (a). We propose utilizing instance-conditional direction to promote the diversity of synthetic features as it can construct customized text description directions for each image rather than a fixed shared text description direction, as illustrated in Fig. 3 (b). To create image-specific text, we use a pre-trained caption model (i.e.,ClipCap [22]) to describe each source image. Then a large language model (i.e.,GPT-3 [2]) combines the image attributes extracted from source descriptions with a base text (e.g. “a photo of a”) to obtain image-specific descriptions. The extracted attributes are consistent between source and target images, only differing in the “domain” description. For a detailed example, refer to Fig. 2, with additional information in Appendix 1.3 explaining the image-specific description construction.
[…]
@@
After obtaining potential visual attributes, we select relevant attributes by ranking their cosine similarity scores to image features within the CLIP joint space.
Subsequently, we train a mapper network with our proposed instance-conditional local adaptation loss $\mathcal{L}_{ld}$ where the optimization objective is defined as aligning the normalized difference between each source image feature $I(x_{s})$ and the adapted feature $F_{m}^{k}(I(x_{s}))$ with the normalized difference between the image-specific target domain description feature $t^{\prime}_{t}$ and the corresponding source domain description feature $t^{\prime}_{s}$ calculated by the text encoder $T$:
@@
Stochastic text feature augmentation: The effectiveness of CLIP-based adaptation loss relies on the assumption that image-text features are well aligned in the joint embedding of CLIP. However, Liang [20] demonstrated that there exists a modality gap in the CLIP embedding that can separate image and text features, impacting cross-modality transferability. We discern this modality gap also impacts adaptation in text-guided image synthesis. Separately, [19] introduced perturbing image features to span the space of domains. Inspired by it, we propose stochastically perturb text features $t^{\prime}_{s}$ and $t^{\prime}_{t}$ to span the text representation space, bringing text features closer to image features before optimizing the adaptation loss $\mathcal{L}_{ld}$ in Eq. 1. Specifically, $t^{\prime}_{s}$ and $t^{\prime}_{t}$ are augmented by adding Gaussian noise $z$ sampled from a standard normal distribution and scaled by a hyperparameter $\beta$, such that $t^{\prime}_{s}=t^{\prime}_{s}+\gamma\cdot z$ and $t^{\prime}_{t}=t^{\prime}_{t}+\gamma\cdot z$. The perturbed text features are then renormalized. We show this stochastic augmentation can close the modality gap and improve image feature augmentation in section 4.3.
@@
3.2 Finetuning CLIP with Synthesized Features
@@
4 Experiments
4.1 Experimental Setup
Implementation details. For PACS, OfficeHome, and DomainNet, we employ a ViT-B/16 based CLIP [29] model as our pre-trained vision-language model, as well as for other baselines. For NICO++, we use a ResNet50-based CLIP model. We build a 2-layer MLP as our mapping network and employ the ClipCap [22] as the captioning model. We report results averaged over 3 runs throughout all experiments. After conducting a grid search on crucial hyperparameters, such as learning rate and weight decay, the optimal configurations are selected based on in-domain validation accuracy for all methods. For prompt learning-based baselines and our methods, we follow the few-shot evaluation protocol in [41, 29], using 16 images per class for training throughout the two stages and evaluating on complete test sets. However, we employ full-shot training in OfficeHome to evaluate our method’s ability under the full-shot setting. All experiments are carried out on two NVIDIA GeForce RTX3090. See Appendix 1.1 for more details about hyper-parameters, training parameters, and GFLOPs.
4.2 Evaluation
Table 1: Top 1 accuracy % on PACS
(ViT-B/16)
@@ 4.2 Evaluation
Table 2: Top 1 accuracy % on OfficeHome (ViT-B/16)

### v2

# Enhancing Vision-Language Models’ Generalization via Diversity-Driven Novel Feature Synthesis

@@ Abstract
Vision-language foundation models like CLIP have shown impressive zero-shot generalization, but finetuning on downstream datasets can cause overfitting and loss of its generalization ability on unseen domains. Although collecting additional data from new domains of interest is possible, this method is often impractical due to the challenges in obtaining annotated data. To address this, we propose a plug-and-play feature synthesis method called LDFS (Language-Guided Diverse Feature Synthesis) to synthesize new domain features and improve existing CLIP fine-tuning strategies. LDFS has three main contributions: 1) To synthesize novel domain features and promote diversity, we propose an instance-conditional feature augmentation strategy based on a text-guided feature augmentation loss. 2) To maintain feature quality after augmenting, we introduce a pairwise regularizer to preserve augmented feature coherence within the CLIP feature space. 3) We propose to use stochastic text feature augmentation to reduce the modality gap and further facilitate the process of text-guided feature synthesis. Extensive experiments show LDFS’s superiority in improving CLIP’s generalization ability on unseen domains without collecting data from those domains. The code will be made publicly available.
Index Terms:
Vision and language, Domain Generalization, Prompt Learning, Data augmentation.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
In order to further adapt CLIP to downstream tasks, some works propose to fine-tune CLIP via either linear probing or learning optimal soft prompts, such as CoOp [45], CoCoOp [44], MaPle [14], and PromptSRC [15], in a process called prompt learning [45, 40, 46, 34] for downstream datasets. Although these fine-tuning methods can improve CLIP’s performance on downstream datasets, a significant issue is that they can also cause the model to overfit to these datasets, resulting in even worse performance than the original zero-shot CLIP on some related datasets with domain-shift, limiting real-world applicability. One method is to collect additional data from the target domains of interest. However, this approach is often impractical due to the challenges in obtaining data from new or rare domains, further compounded by the costs of annotation for effective adaptation.
@@
To address this issue, we aim to develop a feature augmentation method that can synthesize visual features from novel domains by leveraging the knowledge existing in pre-trained CLIP. Our motivation is based on the observation that the domain-shift direction in CLIP’s text embedding can guide image feature augmentation in CLIP’s image embedding, without requiring access to any images from the target domain [17, 5, 8]. However, directly manipulating the text latent space of CLIP to synthesize either novel domain features (e.g., LADS [5]) or images (e.g., CLIPStyler [17]) cannot effectively improve CLIP’s performance on unseen domains. We attribute this to the tendency of these methods to synthesize visual features that lack diversity within the same class (e.g., LADS in Fig. 4) or generate low-quality images with artifacts (e.g., CLIPStyler in Fig. 5). In this paper, we propose LDFS (Language-Guided Diverse Feature Synthesis), which synthesizes diverse and high-quality novel domain visual features given source domain training data along with text descriptions of the training data and the target domain of interest only, as illustrated in the left of Fig. 1. We achieve high diversity and high-quality synthesized features using three approaches. 1)Promoting feature diversity with instance-conditional augmentation: We notice that existing text-guided feature augmentation and styler transfer methods [5, 8, 17] use a shared, global augmentation direction for all images in each category, leading to homogeneous synthesized features or images. We address this by introducing a novel adaptation loss that utilizes more nuanced, instance-conditional text descriptions instead of uniform descriptions. This enables our method to synthesize more diverse visual features. As illustrated in Fig. 3, rather than applying a global direction from “a photo of a horse” to “a sketch of a horse” for all horse images, our method employs customized descriptions like “a (sketch) photo of a horse with its head raised” or “a (sketch) photo of a horse grazing on a green field” for individual images. This strategy ensures that the augmentation process is informed by the unique attributes of each image, enhancing feature diversity and capturing richer visual semantics. 2) Preserving feature coherence in adapted space: We observe that the feature augmentation process could inadvertently drift synthesized features from the CLIP sphere, negatively impacting the synthesis quality and losing the class information. This occurs because the CLIP-based augmentation loss overlooks the preservation of pairwise cosine distances for adapted features. To mitigate it, we propose a regularizer to explicitly maintain the pairwise relationship for augmented features so that augmented features lie in the CLIP sphere, achieving better synthesis quality. 3) Reducing the cross-modality representation gap: Despite CLIP’s ability to represent text and image features in a joint space, there persists a modality representation gap, as documented in recent studies [20, 11]. We notice that this modality gap weakens the integration of cross-modality features and thus hampers the adaptation process. To address it, we propose adjusting the text representation closer to its image counterpart by introducing controlled perturbations.
@@ II Related Work
Text-guided Image Synthesis. Text-guided image synthesis [8, 17, 25, 16] aims to modify source images through text descriptions. For example, StyleGAN-NADA [8] achieves zero-shot image synthesis by leveraging text prompts, identifying the domain shift in the text space as a directional to guide image generation. Similarly, CLIPStyler [17] finetunes a style transfer network utilizing CLIP embeddings for text-driven style transfer. While these works make progress in generative tasks, they are not designed for tackling domain shift problems and necessitate additional generative models. Moreover, their efficacy is frequently constrained by the quality of the generation process, which often introduce artifacts or wrong classes for downstream adaptation (see Fig. 5). Similarly, LADS [5] and PODA are introduced, proposing to directly manipulate CLIP embeddings to synthesize features for model adaptation. Nevertheless, a common issue among all the above methods is their reliance on a global adaptation direction within the CLIP space, which limits the diversity of synthesized images or features. This sometimes leads to even worse performance than zero-shot CLIP (e.g., LADS compared to CLIP (ZS) in Table II to V and CLIPStyler compared to CLIP (ZS) in Table VII. In contrast to prior efforts, we introduce an instance-conditional local direction adaptation framework to promote both diverse and high-quality visual feature synthesis for effective downstream adaptation.
@@
Model Generalization. Domain adaptation (DA) [9, 12, 37] aims to bridge the distribution gap between the source and target domains so that the model can generalize in target domains. In unsupervised domain adaptation, this is typically achieved by leveraging additional unlabeled data from the target domain and utilizing techniques such as adversarial training [9, 12, 7], and moment matching [27, 21, 3], to minimize distribution disparities, all while reducing the necessity for costly annotations in the target domain. More recently, few-shot domain adaptation [23, 39] has been introduced, allowing models to adapt to a new domain with only a few examples, further relaxing the data requirement. Diverging from DA, our method does not require any data from target domains. Instead, it utilizes descriptive text about the source and target domains, guiding the adaptation process of source images via CLIP [29] models. It’s noteworthy that, while domain generalization [43] also does not require target domain data, it typically requires training on multiple source domain data. Although some methods [33, 28, 36] only require a single source domain data (called single domain generalization), they often lead to limited improvements in data with domain shift. This is in contrast to our method, which is trained with just a single source domain and significantly enhances CLIP’s fintuning performance.
III Method
@@
In this section, we present our framework based on an instance-conditional local adaptation strategy, a pair-wise regularizer, and text feature augmentation, which synthesizes diverse and high-quality target domain features rather than using original features only. To formulate the task, consider a training set $\mathcal{D}_{train}=\{(x_{s},y_{s})\}_{s=1}^{S}$ where $x_{s}$ is an input image from source domain, $y_{s}$ is the corresponding ground truth, and S is the size of the source training dataset. The task also includes the description $t_{s}$ of the source training domain (e.g., “a real photo of a {}”) plus the class name (e.g. horse), and a collection of $k$ descriptions $\{t^{i}_{t}\}_{i=1}^{k}$ of the $k$ unseen target domains $\{\mathcal{D}^{i}_{unseen}\}_{i=1}^{k}$ that cover the same classes as the source dataset. The goal is to adapt the model to work well on the $k$ unseen target domains during testing time.
@@
Our two-stage method, shown in Fig. 2, begins with extracting source image features $f=I(x_{s})$ using CLIP’s frozen image encoder. Then, we enrich both source $t_{s}$ and target text descriptions $t_{t}$ with attributes by querying GPT3. The refined descriptions are encoded by CLIP’s text encoder into $t^{\prime}_{s}$ and $t^{\prime}_{t}$. To bridge the modality gap, we apply stochastic, data-independent noise to span the space of text features. A mapper network then maps the source image features $f$ into diverse target features $f^{\prime}$. The mapping process is guided by the domain gap direction between $t^{\prime}_{s}$ and $t^{\prime}_{t}$ using a local adaptation loss $\mathcal{L}_{ld}$. We also employ a sphere regularizer to constrain the synthetic features $f^{\prime}$ within the CLIP sphere. In stage 2, we combine the source image features $f$ with synthetic features $f^{\prime}$ and finetune them to enhance the model performance in unseen domains.
III-A Text-Guided Diverse Feature Synthesis
@@
Instance-conditional local adaptation loss: In the generative model community, existing text-guided image synthesis methods use CLIP embeddings to guide additional generative models like StyleGAN to translate images through losses like global CLIP loss [25] or directional CLIP loss [8, 5]. However, existing CLIP losses utilize a fixed global direction for all images in a class, limiting the diversity of synthetic images, as illustrated in Fig. 3 (a). We propose utilizing instance-conditional direction to promote the diversity of synthetic features as it can construct customized text description directions for each image rather than a fixed shared text description direction, as illustrated in Fig. 3 (b). To create attributes-level text, we query GPT3 [2] to describe each source image with a semantically relevant attribute for each class through the following question:
@@
Notice that the generated attributes does not necessary describe the corresponding image, but it is used to improve the synthetic feature diversity.
After obtaining potential visual attributes, we select relevant attributes by ranking their cosine similarity scores to image features within the CLIP joint space. Subsequently, we train a mapper network with our proposed instance-conditional local adaptation loss $\mathcal{L}_{ld}$ where the optimization objective is defined as aligning the normalized difference between each source image feature $I(x_{s})$ and the adapted feature $F_{m}^{k}(I(x_{s}))$ with the normalized difference between the image-specific target domain description feature $t^{\prime}_{t}$ and the corresponding source domain description feature $t^{\prime}_{s}$ calculated by the text encoder $T$:
@@
Stochastic text feature augmentation: The effectiveness of CLIP-based adaptation loss relies on the assumption that image-text features are well aligned in the joint embedding of CLIP. However, Liang et al.[20] demonstrated that there exists a modality gap in the CLIP embedding that can separate image and text features, impacting cross-modality transferability. We discern this modality gap also impacts adaptation in text-guided image synthesis. Separately, [19] introduced perturbing image features to span the space of domains. Inspired by it, we propose stochastically perturb text features $t^{\prime}_{s}$ and $t^{\prime}_{t}$ to span the text representation space, bringing text features closer to image features before optimizing the adaptation loss $\mathcal{L}_{ld}$ in Eq. 1. Specifically, $t^{\prime}_{s}$ and $t^{\prime}_{t}$ are augmented by adding Gaussian noise $z$ sampled from a standard normal distribution and scaled by a hyperparameter $\beta$, such that $t^{\prime}_{s}=t^{\prime}_{s}+\gamma\cdot z$ and $t^{\prime}_{t}=t^{\prime}_{t}+\gamma\cdot z$. The perturbed text features are then renormalized. We show this stochastic augmentation can close the modality gap and improve image feature augmentation in section IV-C.
@@
III-B Finetuning CLIP with Synthesized Features
@@
IV Experiments
IV-A Experimental Setup
TABLE I: Hyperparameters for training on the four datasets. LR1 and WD1 are the learning rate and weight decay in stage 1. LR2 and WD2 are the learning rate and weight decay in stage 2. $\alpha$ and $\beta$ are weighting hyperparameters for the total loss in stage 1. $\lambda$ is the noise level for stochastic text feature augmentation.
Datasets LR1 WD1 LR2 WD2 $\alpha$ $\beta$ $\lambda$
PACS [18] 1e-3 5e-2 2e-3 5e-4 0.5 0.5 0.1
OfficeHome [32] 1e-2 5e-2 2e-3 1e-4 0.5 1 0.01
DomainNet [26] 1e-2 5e-2 2e-3 1e-4 0.5 1 0.01
NICO++ [41] 1e-2 5e-2 2e-3 1e-4 0.5 0.75 0.08
TABLE II: Top 1 accuracy % on PACS (ViT-B/16).
@@ IV-A Experimental Setup
Implementation details. For PACS, OfficeHome, and DomainNet, we employ a ViT-B/16 based CLIP [29] model as our pre-trained vision-language model, as well as for other baselines. For NICO++, we use a ResNet50-based CLIP model. We build a 2-layer MLP as our mapping network and employ the ClipCap [22] as the captioning model. We report results averaged over 3 runs throughout all experiments. After conducting a grid search on crucial hyperparameters, such as learning rate and weight decay, the optimal configurations are selected based on in-domain validation accuracy for all methods. For prompt learning-based baselines and our methods, we follow the few-shot evaluation protocol in [45, 29], using 16 images per class for training throughout the two stages and evaluating on complete test sets. However, we employ full-shot training in OfficeHome to evaluate our method’s ability under the full-shot setting. All experiments are carried out on two NVIDIA GeForce RTX3090. The hyper-parameters are displayed in Table I. A for training parameters and GFLOPs, the training parameters for feature augmentation are 0.39M x the number of target domains, which is a very small portion of the parameter of ViT-b/16 based CLIP (91.16M+ 86.29M). GFLOPs (inference) is 187.4, which is comparable its baseline CoOp of 187.2.
IV-B Evaluation
TABLE III: Top 1 accuracy % on OfficeHome (ViT-B/16).